# Deflated Dynamics Value Iteration

### v1

@@ 8.6.  Non-asymptotic convergence analysis of DDTD
$Y^{k+1}=Y^{k}+\eta_{k}(f(Y^{k},\zeta_{k})-Y^{k}+Z_{k})$ (9)
for $k=0,1,\dots$, where $Y^{k}\in^{n}$, $\eta_{k}\in^{+}$, $\{\zeta_{k}\}_{k=0,1,\dots}$ are Markov chain with unique distribution $\mu$ and transition matrix $P$. Let $F(y)=\mathbf{E}[f(y,\zeta_{k})]$ and $\mathcal{F}^{k}=\sigma(Y^{i},M^{i},Z_{i},\eta_{i}1\leq i\leq t)$. Define mixing time as $t_{\delta}=\{\min{k\geq 0,:\max_{y}\|P^{k}(\cdot,y)-\mu(\cdot)\|_{\text{TV}}}<\delta\}$ where $\|\cdot\|_{\text{TV}}$ stands for the total variation distance. Let $Y^{\star}$ be fixed point of $F(y)$. Then following proposition holds.
Proposition 8.3.
[Chen et al., 2023, Theorem 1] Let (i) $\{\zeta_{k}\}_{k=0,1,\dots}$ is aperiodic and irreducible Markov chain, (ii) $F$ is $\beta$-contraction respect to some $\|\cdot\|_{c}$, (iii) $f(\cdot,z):^{n}\rightarrow^{n}$ is $A_{1}$-uniformly Lipschitz with respect to $\|\cdot\|_{c}$ and $\|f(0,y)\|_{c}\leq B_{1}$ for any $y$, (iv) $\{M^{k},\mathcal{F}^{k}\}_{k=0,1,\dots,}$ are martingale difference sequence, and (v) $\mathbf{E}[Z^{k+1}\,|\,\mathcal{F}^{k}]\leq A_{2}+B_{2}\|Y^{k}\|^{2}$ for some $A_{2},B_{2}>0$. Then, if $\eta_{k}=\eta$,
@@
for all $k\geq K$, where $\{\eta_{k}\}_{k=0,1,\dots}$ is nonincreasing sequence satisfying $\sum^{k-1}_{i=k-t_{k}}\alpha_{i}\leq\min\left\{\frac{d_{2}}{d_{3}A^{2}},\frac{1}{4A}\right\}$ for all $k\geq t_{k}$, $K=\min\{k\geq 0:k\geq t_{k}\}$, $A=A_{1}+A_{2}+1$, $B=B_{1}+B_{2}$, $c_{1}=(\|Y^{0}-Y^{\star}\|_{c}+\|Y^{0}\|_{c}+A/B)^{2},c_{2}=(A\|Y^{\star}\|_{c}+B)^{2}$, $l_{cs}\|\cdot\|_{s}\leq\|\cdot\|_{c}\leq u_{cs}\|\cdot\|_{s}$ for chosen norm $\|\cdot\|_{s}$ such that $\frac{1}{2}\|\cdot\|^{2}_{s}$ is $L$-smooth function with respect to $\|\cdot\|_{s}$, $d_{1}=\frac{1+\theta u^{2}_{cs}}{1+\theta l^{2}_{cs}}$, $d_{2}=1-\beta d^{1/2}_{1}$, and $d_{3}=\frac{114L(1+\theta u^{2}_{cs})}{\theta l^{2}_{cs}}$ such that $\theta$ is chosen satisfying $\beta^{2}\leq(d_{1})^{-1}$.
DDTD (4) with $\alpha=1$ is equivalent to
$W^{k+1}=W^{k}+\eta_{k}(f(W^{k},X_{k},X^{\prime}_{k})-W^{k}+Z_{k})$
where $Z_{k}=0$, and $f(W^{k},X,X^{\prime})(x)=\mathbf{1}_{\{X=x\}}(r^{\pi}(x)+\gamma(\left(I-\gamma E_{s}\right)^{-1}W^{k})(X^{\prime})-\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}W^{k}(x)-W^{k}(x))+W^{k}(x)$. As we showed in Section 8.5, $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ is irreducible and aperiodic Markov chain with a invariant measure $\mu$. Furthermore, mixing time of $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ is $1$.
@@
$=\mathbf{E}_{\mu\sim\{X,X^{\prime}\}}[\mathbf{1}_{\{X=x\}}(r^{\pi}(x)+\gamma(\left(I-\gamma E_{s}\right)^{-1}W)(X^{\prime})-\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}W(x)-W(x))+W(x)\|X]$
@@
$=\frac{1}{n}(r^{\pi}(x)+\gamma(\mathcal{P}^{\pi}-E_{s})\left(I-\gamma E_{s}\right)^{-1}W(x))+(1-\frac{1}{n})W(x).$
Let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $\mathcal{P}^{\pi}$ sorted in decreasing order of magnitude with ties broken arbitrarily. Let $A_{\text{DDTD}}=\gamma(\mathcal{P}^{\pi}-E_{s})\left(I-\gamma E_{s}\right)^{-1}$. Since spectral radius is the infimum of matrix norm, for any $\epsilon>0$, there exist $\|\cdot\|_{\epsilon}$ such that $\rho(\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I)\leq\|\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I\|_{\epsilon}\leq\rho(\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I)+\epsilon$ where $\rho(\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I)=\max_{\lambda\in\{\lambda_{s+1},\dots,\lambda_{n}\}}\{|1-\frac{1}{n}+\frac{\gamma}{n}\lambda|\}$.
$\|f(0,X,X^{\prime})\|_{\epsilon}=\|\mathbf{1}_{x=X}r^{\pi}(x)\|_{\epsilon}\leq\|r^{\pi}\|_{\epsilon}=B_{\epsilon}$, and $f(W,X,X^{\prime})(x)-f(W^{\prime},X,X^{\prime})(x)=\mathbf{1}_{\{X=x\}}(\gamma\left(I-\gamma E_{s}\right)^{-1}(W-W^{\prime})(X^{\prime})-(\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}(W-W^{\prime}))(x))$ implies that $\|f(W,X,X^{\prime})-f(W^{\prime},X,X^{\prime})\|_{\epsilon^{\prime}}\leq(\|(\gamma\left(I-\gamma E_{s}\right)^{-1}\|_{\epsilon}+\|(\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}\|_{\epsilon})\|W-W^{\prime}\|_{\epsilon}\leq A_{\epsilon}\|W-W^{\prime}\|_{\epsilon}$.
$F(W)=\mathbf{E}_{\mu\sim\{X,X^{\prime}\}}[f(W,X,X^{\prime})]=\frac{1}{n}(r^{\pi}+\gamma(\mathcal{P}^{\pi}-E_{s})\left(I-\gamma E_{s}\right)^{-1}W)+(1-\frac{1}{n})W$. $\|F(W_{1})-F(W_{2})\|_{\epsilon}=\|(\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I)(W_{1}-W_{2})\|_{\epsilon}\leq(\frac{1}{n}\rho+\epsilon+(1-\frac{1}{n}))\|W_{1}-W_{2}\|_{\epsilon}$ and $F(W)$ has fixed point $W^{\pi}$ since $\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I$ and $A_{\text{DDTD}}$ share same fixed point.
Thus, by Theorem 8.3 and $\mathbf{E}[\|V^{k}-V^{\pi}\|_{\epsilon}^{2}]\leq\|(I-\gamma E_{s})^{-1}\|_{\epsilon}^{2}\mathbf{E}[\|W^{k}-W^{\star}\|_{\epsilon}^{2}]$, we have following non-asymptotic convergence result of DDTD.
Theorem 8.4 (DDTD with constant stepsize).
@@
We note that $a_{\epsilon},b_{\epsilon},c_{\epsilon},d_{\epsilon}$ of Theorem 8.4 can be obtained by plugging $A_{1}=A_{\epsilon},A_{2}=0,B_{1}=B_{\epsilon},B_{2}=0$ in Proposition 8.3.
Theorem 8.5 (DDTD with diminishing stepsize).
Let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $\mathcal{P}^{\pi}$ sorted in decreasing order of magnitude with ties broken arbitrarily. For any $\epsilon>0$, there exist $a_{\epsilon},b_{\epsilon},c_{\epsilon},d_{\epsilon}>0$ and $\|\cdot\|_{\epsilon}$ such that for $\alpha=1$ and $\eta_{k}=\frac{1}{(k+b_{\epsilon})(1-c_{\epsilon}\rho)}$, DDTD exhibits the rate
$\mathbf{E}[\|V^{k}-V^{\pi}\|^{2}]\leq\frac{a_{\epsilon}}{k+b_{\epsilon}}+\left(\frac{1}{1-c_{\epsilon}\rho}\right)^{2}\frac{d_{\epsilon}}{k+b_{\epsilon}}$
@@
Again, we note that $a,b,c,$ of Theorem 8.5 can be obtained by plugging $A_{1}=A_{\epsilon},A_{2}=0,B_{1}=B_{\epsilon},B_{2}=0$ in Proposition 8.3.
9.  DDVI with QR Iteration, AutoPI and AutoQR
9.1.  DDVI with QR Iteration
Algorithm 2 Rank-$s$ DDVI with QR Iteration
Initialize $\alpha$, $s$, and $V^{0}$
$\{\lambda_{i}\}^{s}_{i=1},\{u_{s}\}^{s}_{i=1}$ = QRiteration($\mathcal{P}^{\pi}$, s)
for $k=0,\dots,K-1$ do
The QR iteration of Algorithm 2 can be carried out in an automated manner similar to AutoPI. We refer to this as AutoQR and formally describe the algorithm in Appendix 9.
9.2.  DDVI with AutoPI and AutoQR
@@
is the iterates of a power iteration with respect to $(1-\alpha)(I-\alpha\gamma E_{s})^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1}$. In Section 8.1, we show that
@@
For $\alpha\approx 1$, we expect that spectral radius of matrix is $1-\alpha+\gamma\alpha\lambda_{s+1}$ and $(1-\alpha)I+\alpha\gamma\mathcal{P}^{\pi}+U_{s}D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}$ and $\mathcal{P}^{\pi}+U_{s}D_{s,\frac{(\lambda_{i}\gamma-1)\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}$ have same top eigenvector. With same argument in Section 4.2, we can recover the $s+1$-th eigenvector of $\mathcal{P}^{\pi}$ by [Bru et al., 2012, Proposition 5]. Leveraging this observation, we formalize this approach in Algorithm 3.
@@
is the iterates of a power iteration with respect to $(1-\alpha)(I-\alpha\gamma E_{s})^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1}$. In Section 8.1, we show that
@@
For $\alpha\approx 1$, we expect that spectral radius of matrix is $1-\alpha+\gamma\alpha\lambda_{s+1}$ and $(1-\alpha)I+\alpha\gamma\mathcal{P}^{\pi}+U_{s}R_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}$ and $\mathcal{P}^{\pi}+U_{s}R_{s,\frac{(\lambda_{i}\gamma-1)\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}U^{\top}_{s}$ have same top eigenvector. With same argument in Section 4.2, for large $k$, we expect $\frac{W^{k+1}-W^{k}}{\|W^{k+1}-W^{k}\|}\approx w,$ where $w$ is the top eigenvector of $\mathcal{P}^{\pi}-U_{s}R_{s,\frac{(\lambda_{i}\gamma-1)\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}U^{\top}_{s}$. Thus, by orthornormalizing $w$ against $u_{1},\dots,u_{s}$, we can obtain top $s+1$-th Schur vector of $\mathcal{P}^{\pi}$ Saad [2011, Section 4.2.4]. Leveraging this observation, we propose DDVI with Automatic QR Iteration (AutoQR) and formalize this in Algorithm 4.
@@
10.  Environments

### v2

@@ B.6 Non-asymptotic convergence analysis of DDTD
$Y^{k+1}=Y^{k}+\eta_{k}(X_{k})(f(Y^{k},\zeta_{k})-Y^{k}+Z_{k})$ (15)
for $k=0,1,\dots$, where $Y^{k}\in^{n}$, $\eta_{k}(X_{k})\in^{+}$, $\{\zeta_{k}\}_{k=0,1,\dots}$ are Markov chain with unique distribution $\mu$ and transition matrix $P$. Let $F(y)=\mathbf{E}[f(y,\zeta_{k})]$ and $\mathcal{F}^{k}=\sigma(Y^{i},M^{i},Z_{i},\eta_{i}1\leq i\leq t)$. Define mixing time as $t_{\delta}=\{\min{k\geq 0,:\max_{y}\|P^{k}(\cdot,y)-\mu(\cdot)\|_{\text{TV}}}<\delta\}$ where $\|\cdot\|_{\text{TV}}$ stands for the total variation distance. Let $Y^{\star}$ be fixed point of $F(y)$. Then following proposition holds.
Proposition B.3.
(Chen et al., 2023, Theorem 1) Let (i) $\{\zeta_{k}\}_{k=0,1,\dots}$ is aperiodic and irreducible Markov chain, (ii) $F$ is $\beta$-contraction respect to some $\|\cdot\|_{c}$, (iii) $f(\cdot,z):^{n}\rightarrow^{n}$ is $A_{1}$-uniformly Lipschitz with respect to $\|\cdot\|_{c}$ and $\|f(0,y)\|_{c}\leq B_{1}$ for any $y$, (iv) $\{M^{k},\mathcal{F}^{k}\}_{k=0,1,\dots,}$ are martingale difference sequence, and (v) $\mathbf{E}[Z^{k+1}\,|\,\mathcal{F}^{k}]\leq A_{2}+B_{2}\|Y^{k}\|^{2}$ for some $A_{2},B_{2}>0$. Then, if $\eta_{k}=\eta$,
@@
for all $k\geq K$, where $\{\eta_{k}(X_{k})\}_{k=0,1,\dots}$ is nonincreasing sequence satisfying $\sum^{k-1}_{i=k-t_{k}}\alpha_{i}\leq\min\left\{\frac{d_{2}}{d_{3}A^{2}},\frac{1}{4A}\right\}$ for all $k\geq t_{k}$, $K=\min\{k\geq 0:k\geq t_{k}\}$, $A=A_{1}+A_{2}+1$, $B=B_{1}+B_{2}$, $c_{1}=(\|Y^{0}-Y^{\star}\|_{c}+\|Y^{0}\|_{c}+A/B)^{2},c_{2}=(A\|Y^{\star}\|_{c}+B)^{2}$, $l_{cs}\|\cdot\|_{s}\leq\|\cdot\|_{c}\leq u_{cs}\|\cdot\|_{s}$ for chosen norm $\|\cdot\|_{s}$ such that $\frac{1}{2}\|\cdot\|^{2}_{s}$ is $L$-smooth function with respect to $\|\cdot\|_{s}$, $d_{1}=\frac{1+\theta u^{2}_{cs}}{1+\theta l^{2}_{cs}}$, $d_{2}=1-\beta d^{1/2}_{1}$, and $d_{3}=\frac{114L(1+\theta u^{2}_{cs})}{\theta l^{2}_{cs}}$ such that $\theta$ is chosen satisfying $\beta^{2}\leq(d_{1})^{-1}$.
DDTD in equation 8 with $\alpha=1$ is equivalent to
$W^{k+1}=W^{k}+\eta_{k}(X_{k})(f(W^{k},X_{k},X^{\prime}_{k})-W^{k}+Z_{k}),$
where $Z_{k}=0$, and $f(W^{k},X,X^{\prime})(x)=\mathbf{1}_{\{X=x\}}(r^{\pi}(x)+\gamma(\left(I-\gamma E_{s}\right)^{-1}W^{k})(X^{\prime})-\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}W^{k}(x)-W^{k}(x))+W^{k}(x)$. As we showed in Section B.5, $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ is irreducible and aperiodic Markov chain with an invariant measure $\mu$. Furthermore, mixing time of $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ is $1$.
We also have
@@
$=\mathbf{E}_{\mu\sim\{X,X^{\prime}\}}[\mathbf{1}_{\{X=x\}}(r^{\pi}(x)+\gamma(\left(I-\gamma E_{s}\right)^{-1}W)(X^{\prime})-\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}W(x)-W(x))+W(x)\mid X]$
@@
$=\frac{1}{n}(r^{\pi}(x)+\gamma(\mathcal{P}^{\pi}-E_{s})\left(I-\gamma E_{s}\right)^{-1}W(x))+(1-\frac{1}{n})W(x).$ (16)
Let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $\mathcal{P}^{\pi}$ sorted in the decreasing order of magnitude with ties broken arbitrarily. Let $A_{\text{DDTD}}=\gamma(\mathcal{P}^{\pi}-E_{s})\left(I-\gamma E_{s}\right)^{-1}$. Since spectral radius is the infimum of matrix norm, for any $\epsilon>0$, there exists $\|\cdot\|_{\epsilon}$ such that $\rho(\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I)\leq\|\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I\|_{\epsilon}\leq\rho(\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I)+\epsilon$ where $\rho(\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I)=\max_{\lambda\in\{\lambda_{s+1},\dots,\lambda_{n}\}}\{|1-\frac{1}{n}+\frac{\gamma}{n}\lambda|\}$.
The upper bound $\|f(0,X,X^{\prime})\|_{\epsilon}=\|\mathbf{1}_{x=X}r^{\pi}(x)\|_{\epsilon}\leq\|r^{\pi}\|_{\epsilon}=B_{\epsilon}$ and $f(W,X,X^{\prime})(x)-f(W^{\prime},X,X^{\prime})(x)=\mathbf{1}_{\{X=x\}}(\gamma\left(I-\gamma E_{s}\right)^{-1}(W-W^{\prime})(X^{\prime})-(\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}(W-W^{\prime}))(x))$ implies that
$\|f(W,X,X^{\prime})-f(W^{\prime},X,X^{\prime})\|_{\epsilon^{\prime}}$ $\leq\left(\|(\gamma\left(I-\gamma E_{s}\right)^{-1}\|_{\epsilon}+\|(\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}\|_{\epsilon}\right)\|W-W^{\prime}\|_{\epsilon}$
$\leq A_{\epsilon}\|W-W^{\prime}\|_{\epsilon}.$
As shown in equation B.6, we have $F(W)=\mathbf{E}_{\mu\sim\{X,X^{\prime}\}}[f(W,X,X^{\prime})]=\frac{1}{n}(r^{\pi}+\gamma(\mathcal{P}^{\pi}-E_{s})\left(I-\gamma E_{s}\right)^{-1}W)+(1-\frac{1}{n})W$. This shows that
$\left\|F(W_{1})-F(W_{2})\right\|_{\epsilon}=\left\|\left(\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I\right)(W_{1}-W_{2})\right\|_{\epsilon}\leq\left(\frac{1}{n}\rho+\epsilon+(1-\frac{1}{n})\right)\left\|W_{1}-W_{2}\right\|_{\epsilon}$
and $F(W)$ has fixed point $W^{\pi}$ since $\frac{1}{n}A_{\text{DDTD}}+(1-\frac{1}{n})I$ and $A_{\text{DDTD}}$ share same fixed point.
Thus, by Theorem B.3 and $\mathbf{E}[\|V^{k}-V^{\pi}\|_{\epsilon}^{2}]\leq\|(I-\gamma E_{s})^{-1}\|_{\epsilon}^{2}\mathbf{E}[\|W^{k}-W^{\star}\|_{\epsilon}^{2}]$, we have the following two non-asymptotic convergence results for DDTD – one for constant stepsize and the other for a diminishing stepsize.
Theorem B.4 (DDTD with constant stepsize).
@@
We note that $a_{\epsilon},b_{\epsilon},c_{\epsilon},d_{\epsilon}$ of Theorem B.4 can be obtained by plugging $A_{1}=A_{\epsilon},A_{2}=0,B_{1}=B_{\epsilon},B_{2}=0$ in Proposition B.3.
Theorem B.5 (DDTD with diminishing stepsize).
Let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $\mathcal{P}^{\pi}$ sorted in decreasing order of magnitude with ties broken arbitrarily. For any $\epsilon>0$, there exist $a_{\epsilon},b_{\epsilon},c_{\epsilon},d_{\epsilon}>0$ and $\|\cdot\|_{\epsilon}$ such that for $\alpha=1$ and $\eta_{k}(X_{k})=\frac{1}{(k+b_{\epsilon})(1-c_{\epsilon}\rho)}$, DDTD exhibits the rate
$\mathbf{E}[\|V^{k}-V^{\pi}\|_{\epsilon}^{2}]\leq\frac{a_{\epsilon}}{k+b_{\epsilon}}+\left(\frac{1}{1-c_{\epsilon}\rho}\right)^{2}\frac{d_{\epsilon}}{k+b_{\epsilon}}$
@@
Again, we note that $a,b,c,$ of Theorem B.5 can be obtained by plugging $A_{1}=A_{\epsilon},A_{2}=0,B_{1}=B_{\epsilon},B_{2}=0$ in Proposition B.3.
Appendix C DDVI with QR Iteration, AutoPI and AutoQR
C.1 DDVI with AutoPI and AutoQR
@@
are the iterates of a power iteration with respect to $(1-\alpha)(I-\alpha\gamma E_{s})^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1}$. In Section B.1, we show that
@@
For $\alpha\approx 1$, we expect that spectral radius of matrix be $1-\alpha+\gamma\alpha\lambda_{s+1}$ and $(1-\alpha)I+\alpha\gamma\mathcal{P}^{\pi}+U_{s}D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}$ and $\mathcal{P}^{\pi}+U_{s}D_{s,\frac{(\lambda_{i}\gamma-1)\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}$ have the same top eigenvector. With the same argument as in Section 4.2, we can recover the $s+1$-th eigenvector of $\mathcal{P}^{\pi}$ by Bru et al. (2012, Proposition 5). Leveraging this observation, we formalize this approach in Algorithm 4.
@@
is the iterates of a power iteration with respect to $(1-\alpha)(I-\alpha\gamma E_{s})^{-1}+\alpha\gamma(\mathcal{P}^{\pi}-E_{s})(I-\alpha\gamma E_{s})^{-1}$. In Section B.1, we show that
@@
For $\alpha\approx 1$, we expect that spectral radius of matrix is $1-\alpha+\gamma\alpha\lambda_{s+1}$ and $(1-\alpha)I+\alpha\gamma\mathcal{P}^{\pi}+U_{s}R_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}V^{\top}_{s}$ and $\mathcal{P}^{\pi}+U_{s}R_{s,\frac{(\lambda_{i}\gamma-1)\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}U^{\top}_{s}$ have the same top eigenvector. With the same argument as in Section 4.2, for large $k$, we expect $\frac{W^{k+1}-W^{k}}{\|W^{k+1}-W^{k}\|}\approx w,$ where $w$ is the top eigenvector of $\mathcal{P}^{\pi}-U_{s}R_{s,\frac{(\lambda_{i}\gamma-1)\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}U^{\top}_{s}$. Thus, by orthornormalizing $w$ against $u_{1},\dots,u_{s}$, we can obtain top $s+1$-th Schur vector of $\mathcal{P}^{\pi}$ Saad (2011, Section 4.2.4). Leveraging this observation, we propose DDVI with Automatic QR Iteration (AutoQR) and formalize this in Algorithm 5.
@@
Appendix D Environments
Figure 4: Cliffwalk environment. Dark shaded region is goal state with a positive reward. Gray shaded regions are terminal states with negative reward. Arrows show the optimal policy.